OpenStudy (anonymous):

Four pine seedling and 3 fir seedlings cost 29.00. Two pine seedlings and 7 fir seedlings cost $42.00 Determine the cost of each type of seedling.

hero (hero):

4p + 3f = 29 2p + 7f = 42 Multiply the second equation by -2 to get: 4p + 3f = 29 -4p -14f = -84 Now, you are authorized to add. (You can only add when eliminating a variable): -11f = 55 11f = -55 f = -55/11 f = -5 Now find p

OpenStudy (anonymous):

But why do we mulitipy it by negative 2?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

?

hero (hero):

Because you need to do that in order to eliminate a variable